OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center (0, 8), radius = 8

OpenStudy (anonymous):

x^2+(y-8)^2=8^2

OpenStudy (anonymous):

(x)^2 + (y-8)^2=64